Christopher is 2 times as old as Gabriela. Nine years ago, Christopher was 5 times as old as Gabriela.
How old is Christopher now?

Answer by harpazo(655)


Let 2x = Christopher's age
Let x = Gabriela's age
2x - 9 = 5(x-9)
2x - 9 = 5x-45
2x- 5x = -45+9
-3x = -36
x = -36/-3
x = 12
Christopher 2x years old today or 2 times 12 which is 24 years old.